# Self-assembled filament layers in drying sessile droplets: from morphology to electrical conductivity

### v1

@@ II.4 Problem definition and assumptions
We investigate how filament length, concentration, stiffness, and evaporation modes influence deposition patterns. Simulations are performed on a lattice of size $96\times 96\times 48$ nodes, initialized with an equilibrated hemispherical droplet of base radius $h_{0}=a=38$ nodes, as shown in Fig. 1. The evaporation rate is chosen so that complete evaporation for single droplets occurs over $N_{t}=200{,}000$ simulation steps, defining a characteristic interface velocity $v_{c}=a/N_{t}=1.9\times 10^{-4}$. The surface tension $\sigma=2\times 10^{-3}$ is chosen as high as possible so that the droplet has the shape of a spherical cap, but small enough to minimize spurious currents without inducing numerical instability in the filament acceleration integration. We implement a stick-slide evaporation mode: the contact line remains pinned until the contact angle decreases to approximately $40^{\circ}$, after which retraction occurs. The simulation operates in a capillary-dominated regime where gravity is negligible, justified by a small Bond number ($\mathrm{Bo}\ll 1$). We assume isothermal conditions with surfactant-free droplets; consequently, Marangoni flows are neglected ($\mathrm{Ma}\ll 1$[Greyson2017, Hu2006]. The flow is characterized by low Reynolds numbers ($\mathrm{Re}\ll 1$), indicating that viscous forces dampen inertial effects. Solute transport is convection-dominated ($\mathrm{Pe}\gg 1$), with Brownian diffusion suppressed by chain connectivity and transient entanglements [KolahgarAzari2022]. To ensure stable fluid-filament coupling, we utilize a bead mass of $m=60$, resulting in a low Stokes number ($\mathrm{St}\ll 1$). This confirms that the inertial response of the beads is negligible compared to viscous damping. Consequently, the beads effectively trace the local fluid velocity, subject only to confinement by the receding interface and interactions with other beads and the substrate.
@@ III.1 Evaporation of a pure droplet
We first examine the evaporation of a pure droplet and validate the evaporation model by comparing the time evolution of the droplet height $h(r,t)$ and the radial capillary flow $v(r,t)$ with theoretical predictions. The so-called rush-hour behavior, characterized by an increase in $v(r,t)$ over time, is expected in the CCR mode [Marn2011], but is partially suppressed here due to deviations from the ideal spherical-cap shape and dynamic pinning [Li2016, Zhang2015]. Once the receding contact angle of $\theta\approx 40^{\circ}$ is reached, the capillary flow driven by pinning diminishes in the CCA mode [Greyson2017]. Our simulations presented below extend previously published benchmarks that were limited to small contact angles in the CCR mode and uniform evaporation [Xie2018, Nath2025].
Figure 1: Geometry of a spherical cap droplet with radius of curvature $R$, base radius $a$, height $h_{0}$, and local height $h$ at radial position $r$. The droplet rests on a patterned substrate with variable wettability: a circular hydrophilic area (red, $\theta\approx 40^{\circ}$) is located at the center, surrounded by a neutral wetting area (green, $\theta=90^{\circ}$).
@@
which, for small contact angles, simplifies to [Marn2011_2]
@@ III.1.1 Mass conservation
Mass conservation dictates that for perfect contact line pinning, the rate of change of liquid height in an infinitesimal annular element at radius $r$ equals the net radial flux into the element minus the evaporated mass [Deegan2000, Marn2011_2]:
$\frac{\partial h}{\partial t}=-\frac{1}{r}\frac{\partial(rhv)}{\partial r}-\frac{1}{\rho}J\sqrt{1+\left(\frac{\partial h}{\partial r}\right)^{2}}$ (38)
Here, $Q=rhv$ is the radial volume flux, $\rho$ is the liquid density, and $J$ the local evaporative flux. The geometrical factor $\sqrt{1+(\partial_{r}h)^{2}}$ accounts for the local surface inclination.
@@
The time derivative of the local droplet height $h(r,t)$ follows from the chain rule as
@@
with
$\frac{\partial h}{\partial a}=\frac{2h_{0}\left(r^{2}+h_{0}^{2}\right)a}{\left(a^{2}+h_{0}^{2}\right)^{2}},$ (43)
$\frac{\partial h}{\partial h_{0}}=\frac{\left(r^{2}-a^{2}\right)\left(h_{0}^{2}-a^{2}\right)}{\left(h_{0}^{2}+a^{2}\right)^{2}}.$ (44)
This contribution is negligible under perfect pinning ($\frac{\partial a}{\partial t}=0$) but dominates at small contact angles during depinning.
@@ III.1.2 Droplet Height
In the reaction-limited regime ($K=10$) with nearly uniform evaporation, an analytical expression can be obtained. For a droplet in the CCR mode, the height is given by [Nath2025]
@@
where $\langle J\rangle$ represents the spatially averaged flux. The total evaporation time is defined as $t_{e}=h_{0}\rho_{0}/(2\langle J\rangle)$. For $J_{0}=0.0005$, the slope from linear regression agrees with the theoretical prediction within $0.3\%$, confirming that the model reproduces the expected linear decay of the droplet during the CCR stage. Adjusting the threshold parameter controlling the interface width may further improve agreement [Nath2025].
In the diffusion-limited regime ($K\approx 0)$, the non-uniform flux requires numerical integration of $J(r)$ to compute the height evolution. For small droplets at late stages ($h\lesssim 10$ lattice units), deviations appear due to the CGLB discretization limits, with the theory predicting a slightly faster evaporation. This discrepancy is negligible for our analysis, as inertia is minimal and most particles are deposited by this stage. Applying a late-time linear fit to extract a correction factor $S$ (see Eq. 15) restores consistency with our simulations.
@@
III.1.3 Radial Velocity
The particle transport toward the droplet edge during evaporation, known as the coffee-ring effect, is driven by the radial velocity $v(r,t)$. Following Deegan et al. [Deegan2000], it can be expressed as
where $h(r,t)$ is the droplet interface and $J(r,t)$ the local evaporative flux. In the original formulation of Deegan [Deegan2000], the last term $\zeta_{r}$ is absent; here it is introduced as a correction to account for contact line motion (see Eq. 42). The relative radial velocity is defined as $\tilde{v}(r,t)=v(r,t)/v_{c}$. For reaction-limited evaporation with $K\in\mathcal{O}(10)$, contact line pinning drives a radial flow since the evaporation is nearly uniform [MURISIC2011, Larsson2023].
Eq. 47 without the correction term $\zeta_{r}$ assumes perfect contact line pinning. In contrast, in the stick-slide regime, the base radius remains nearly constant during the CCR regime, followed by a pronounced reduction in the CCA mode. Fig 3 benchmarks the CGLB evaporation model against theoretical predictions with and without the correction term, considering the effect of a dynamic contact line on the radial velocity. Both the reaction-limited ($K=10$, blue) and the diffusion-limited ($K\approx 0$, green) regimes show good agreement. For analysis, we exclude the initial portion corresponding to the system’s equilibration. Dashed lines denote the uncorrected theory, while solid lines include the $\zeta_{r}$ correction. Shaded regions indicate uncertainties of $\pm 0.5$ lattice units in the interface position and $1\%$ in the correction factor $S$ used in the CGLB formulation [Nath2025].
@@ III.2 Evaporation of a filament-laden droplet
We now investigate the self-organization of filaments within an evaporating droplet. Fig. 4 visualizes a filament-filled droplet drying on a substrate under reaction-limited and diffusion-limited evaporation regimes. We initialize non-overlapping straight filaments, randomly positioned and oriented within the droplet volume (Fig. 4a and 4e). During evaporation, filaments are first gradually advected toward the contact line (Fig. 4b and 4f). Bending arises primarily from solvation forces at the interface, which cause free chain segments to align with the local curvature of the droplet surface. Additional shape changes result from interparticle interactions and adhesion between deposited segments and the substrate. In this stage, filament segments with sufficient substrate friction overcome solvation forces and remain deposited, while unconstrained segments are carried inward by the receding interface (Fig. 4c and 4g). Near the contact line, filaments are effectively immobilized by friction between the filament and the substrate. Away from the contact line, if one end of a filament is fixed, interparticle forces transmit this constraint along the chain, promoting radial alignment. The filaments finally settle onto the substrate (Fig. 4d and 4h[Park2024], leading to a coffee-ring pattern in the diffusion-limited regime ($K\approx 0$) and a homogeneous deposit in the reaction-limited regime ($K=10$).
@@ III.2.1 Order of Filament Deposits
where $r_{i}$ is the bin’s radial position, $N_{i}$ the number of segments within the bin, $\mathbf{u}_{j}$ is the orientation vector of the $j$-th segment, and $\hat{\mathbf{t}}_{j}$ the local tangent unit vector at the segment midpoint $\mathbf{r}_{j}=(x_{j},y_{j})$:
@@
Here, $n(r)=N(r)/\Delta A(r)$ is the number density within the $i$-th annular bin $r_{i},r_{i+1}]$, where $N(r)$ is the number of beads located within the bin, and $\Delta A(r)=\pi(r_{i+1}^{2}-r_{i}^{2})$ is the area of the annulus. The average density is $\rho_{0}=N_{\text{total}}/(\pi R^{2})$, where $N_{\text{total}}$ is the total number of beads. Values of $g(r)>1$ indicate enrichment, while $g(r)<1$ signifies depletion at the corresponding radial distance.
@@
The tangential nematic order parameter $S_{t}(r)$ and the radial distribution function $g(r)$ capture how confinement, filament length $L$, concentration $c_{v}$, substrate interactions, and dynamic effects - such as contact line pinning and evaporation-driven flow - govern structural organization during droplet evaporation. Fig. 5 presents $S_{t}(r)$ and $g(r)$ for four representative cases obtained by varying the filament length $L$ and the filament concentration $c_{v}$. To improve statistical significance, data were averaged over parameter ranges defined as “small filaments” ($L_{\mathrm{min}}\leq L<L_{\mathrm{small}}$) and “large filaments” ($L_{\mathrm{large}}<L\leq L_{\mathrm{max}}$), as well as “low concentrations” ($c_{\mathrm{min}}\leq c_{v}<c_{\mathrm{low}}$) and “high concentrations” ($c_{\mathrm{high}}<c_{v}\leq c_{\mathrm{max}}$). The corresponding thresholds were $L_{\mathrm{small}}=11$, $L_{\mathrm{large}}=23$, $L_{\mathrm{max}}=31$, $c_{\mathrm{low}}=0.0035$, $c_{\mathrm{high}}=0.0107$, and $c_{\mathrm{max}}\approx 0.0131$. The averaging was performed with step sizes of $\Delta L=2$ and $\Delta c_{v}=0.0012$. Three characteristic radial regions can be distinguished. As shown in Fig. 5, in the outer region ($\tilde{r}\gtrsim 0.75$), where $\tilde{r}=r/a_{max}$ and $a_{max}=32.79$ denotes the average maximal deposit radius, we find $S_{t}(r)>0$, indicating pronounced tangential alignment (corresponding to Fig. 4d and 4h). This alignment results from outward capillary flows during the CCR phase, with filaments advected to the interface adapting to its curvature. In the intermediate region ($0.25\lesssim\tilde{r}<0.75$), $S_{t}(r)<0$ reflects radial alignment. Here, shear gradients during the CCR phase stretch filaments radially, while in the CCA phase the receding contact line reinforces this orientation as anchored segments resist the inward motion of free segments. In the inner region ($\tilde{r}\lesssim 0.25$), high particle density or low radial flow and excluded-volume effects suppress radial alignment, leading to weak tangential ordering as filaments aggregate into an approximately isotropic central cluster.
The magnitude and spatial extent of these alignment patterns depend on filament length: longer filaments show stronger tangential alignment, particularly outside the inner region, due to fewer inter-particle interactions, whereas shorter filaments experience stronger interactions that disrupt orientational order (top row of Fig. 5).
Evaporation conditions further modulate this behavior. In the diffusion-limited regime ($K\approx 0$), strong capillary flows persist into the CCA phase (see Fig. 3), transporting particles toward the contact line and enhancing tangential alignment (see Fig. 5). In contrast, in the reaction-limited regime ($K=10$), uniform evaporation suppresses capillary flows and inward contact line motion dominates, resulting in weaker tangential order. For short filaments at low concentrations, this difference is negligible due to reduced inter-particle interactions.
The radial density profile $g(r)$ supports these trends. For $K=10$, central enrichment and edge depletion are observed: short filaments exhibit a weak coffee-ring peak near $\tilde{r}\approx 0.85$, whereas long filaments remain concentrated near the center and show no distinct coffee-ring peak. Under $K\approx 0$, central depletion and edge accumulation occur, yet long filaments still exhibit a comparatively uniform distribution.

### v2

@@ II.4 Problem definition and assumptions
We investigate how filament length, concentration, stiffness, and evaporation modes influence deposition patterns. Simulations are performed on a lattice of size $96\times 96\times 48$ nodes, initialized with an equilibrated hemispherical droplet of base radius $h_{0}=a=38$ nodes corresponding to a droplet radius of approximately $2\mu m$ [14], as shown in Fig. 1. This scale represents water with dynamic viscosity $\eta_{w}=1\times 10^{-3}\,\mathrm{Pa\cdot s}$, density $\rho_{w}=10^{3}\,\mathrm{kg/m^{3}}$, and surface tension $\sigma_{w}=7.2\times 10^{-2}\,\mathrm{N/m}$. Filaments have a width of $\approx 50\,\mathrm{nm}$ and a maximum length of $\approx 1.5\,\mu m$. The evaporation rate is chosen so that complete evaporation for single droplets occurs over $N_{t}=200{,}000$ simulation steps, defining a characteristic interface velocity $v_{c}=a/N_{t}=1.9\times 10^{-4}$. The surface tension $\sigma=2\times 10^{-3}$ is chosen as high as possible so that the droplet has the shape of a spherical cap, but small enough to minimize spurious currents without inducing numerical instability in the filament acceleration integration. We implement a stick-slide evaporation mode: the contact line remains pinned until the contact angle decreases to approximately $40^{\circ}$, after which retraction occurs [50]. The evaporation flux scale $J(0)$ is chosen to reflect realistic printing conditions, ensuring that key dimensionless numbers in the simulations match experimental values. The simulation operates in a capillary-dominated regime where gravity is negligible, justified by a small Bond number ($\mathrm{Bo}\ll 1$). We assume isothermal conditions with surfactant-free droplets; consequently, Marangoni flows are neglected ($\mathrm{Ma}\ll 1$[22, 28]. The flow is characterized by low Reynolds numbers ($\mathrm{Re}\ll 1$), indicating that viscous forces dampen inertial effects. Solute transport is convection-dominated ($\mathrm{Pe}\gg 1$), with Brownian diffusion suppressed by chain connectivity [68, 64]. To ensure stable fluid-filament coupling, we utilize a bead mass of $m=60$, resulting in a low Stokes number ($\mathrm{St}\ll 1$). This confirms that the inertial response of the beads is negligible compared to viscous damping. Consequently, the beads effectively trace the local fluid velocity, subject only to confinement by the receding interface and interactions with other beads and the substrate. This is consistent with overdamped filament dynamics in typical printing scenarios.
@@ III.1 Evaporation of a pure droplet
We first examine the evaporation of a pure droplet and validate the evaporation model by comparing the time evolution of the droplet height $h(r,t)$ and the radial capillary flow $v(r,t)$ with theoretical predictions. The so-called rush-hour behavior, characterized by an increase in $v(r,t)$ over time, is expected in the CCR mode [46], but is partially suppressed here due to deviations from the ideal spherical-cap shape and dynamic pinning [42, 74]. Once the receding contact angle of $\theta\approx 40^{\circ}$ is reached, the capillary flow driven by pinning diminishes in the CCA mode [22]. Our simulations presented below extend previously published benchmarks that were limited to small contact angles in the CCR mode and uniform evaporation [71, 49]. While the current study focuses on a dilute filament concentration, we note that the presence of filaments increases the effective viscosity of the suspension [32] and correspondingly suppresses capillary flows.
Figure 1: A resting droplet on a patterned substrate with variable wettability: a central circular hydrophilic region (red, $\theta\approx 40^{\circ}$) surrounded by a neutral wetting area (green, $\theta=90^{\circ}$). (a) Geometry of a spherical cap droplet with radius of curvature $R$, base radius $a$, height $h_{0}$, and local height $h$ at radial position $r$. (b) Stick-slip dynamics illustrated with later time steps shown in darker colors: the dashed line marks the constant contact radius mode (CCR), followed by dotted lines indicating the constant contact angle mode (CCA).
@@
which, for small contact angles, simplifies to [45]
@@ III.1.1 Mass conservation
Mass conservation dictates that for perfect contact line pinning, the rate of change of liquid height in an infinitesimal annular element at radius $r$ equals the net radial flux into the element minus the evaporated mass [16, 45]:
$\frac{\partial h}{\partial t}=-\frac{1}{r}\frac{\partial{\color[rgb]{0,0,0}Q}}{\partial r}-\frac{1}{\rho}J\sqrt{1+\left(\frac{\partial h}{\partial r}\right)^{2}}$ (38)
Here, $Q=rhv$ is the radial volume flux, $\rho$ is the liquid density, $J$ the local evaporative flux. The geometrical factor $\sqrt{1+(\partial_{r}h)^{2}}$ accounts for the local surface inclination.
@@
The time derivative of the local droplet height $h(r,t)$ follows from the chain rule and Eq. 37 as
@@
This contribution is negligible under perfect pinning ($\frac{\partial a}{\partial t}=0$) but dominates at small contact angles during depinning. Including this correction term, the governing equation becomes
$\color[rgb]{0,0,0}\frac{\partial h}{\partial t}=-\frac{1}{r}\frac{\partial{\color[rgb]{0,0,0}Q}}{\partial r}-\frac{1}{\rho}J\sqrt{1+\left(\frac{\partial h}{\partial r}\right)^{2}}-\zeta_{r}.$} (43)
@@ III.1.2 Droplet Height
In the reaction-limited regime ($K=10$) with nearly uniform evaporation, an analytical expression can be obtained. For a droplet in the CCR mode, the height is given by [49]
@@
where $\langle J\rangle$ represents the spatially averaged flux. Here, we use $h(t)$ to denote $h_{r=0}(t)$. The total evaporation time is defined as $t_{e}=h_{0}\rho/(2\langle J\rangle)$. For $J_{0}=0.0005$, the slope obtained from linear regression differs from the theoretical prediction by less than $0.3\%$, confirming that the model reproduces the expected linear decay of the droplet during the CCR stage. Adjusting the threshold parameter controlling the interface width may further improve the agreement [49].
In the diffusion-limited regime ($K\approx 0)$, the non-uniform flux requires numerical integration of $J(r)$ to compute the height evolution. For small droplets at late stages ($h\lesssim 10$ lattice units), deviations appear due to the CGLB discretization limits (the interface is diffusive with a typical thickness of around 6 lattice units), with the theory predicting a slightly faster evaporation. This discrepancy is negligible for our analysis, as inertia is minimal and most particles are deposited by this stage. Applying a late-time linear fit to extract a correction factor $S$ (see Eq. 15) restores consistency with our simulations. Due to the interface thickness, we neglect detailed filament deposition and growth dynamics at the contact line. For droplets with a radius comparable to or smaller than the filament length, this approximation holds well; however, for larger droplet-to-filament ratios, the spatial distribution of deposits and alignment patterns may be significantly affected, since they are dominated by capillary flows.
III.1.3 Radial Velocity
The particle transport toward the droplet edge during evaporation, known as the coffee-ring effect, is driven by the radial velocity $v(r,t)$. Following Deegan et al. [16], it can be expressed as
where $h(r,t)$ is the droplet interface and $J(r,t)$ the local evaporative flux. In the original formulation of Deegan [16], the last term $\zeta_{r}$ is absent; here it is introduced as a correction to account for contact line motion (see Eq. 42). The relative radial velocity is defined as $\tilde{v}(r,t)=v(r,t)/v_{c}$. For reaction-limited evaporation with $K\in\mathcal{O}(10)$, contact line pinning drives a radial flow since the evaporation is nearly uniform [48, 37].
@@ III.1.3 Radial Velocity
Eq. 46 without the correction term $\zeta_{r}$ assumes perfect contact line pinning. In contrast, in the stick-slide regime, the base radius remains nearly constant during the CCR regime, followed by a pronounced reduction in the CCA mode. Fig 3 benchmarks the CGLB evaporation model against theoretical predictions with and without the correction term, considering the effect of a dynamic contact line on the radial velocity. Both the reaction-limited ($K=10$, blue) and the diffusion-limited ($K\approx 0$, green) regimes show good agreement. For analysis, we exclude the initial portion corresponding to the system’s equilibration. Dashed lines denote the uncorrected theory, while solid lines include the $\zeta_{r}$ correction. Shaded regions indicate uncertainties of $\pm 0.5$ lattice units in the interface position and $1\%$ in the correction factor $S$ used in the CGLB formulation [49].
@@ III.2 Evaporation of a filament-laden droplet
We now investigate the self-organization of filaments within an evaporating droplet. Fig. 4 visualizes the dynamic evolution of a filament-filled droplet drying on a substrate under reaction-limited and diffusion-limited evaporation regimes. We initialize non-overlapping straight filaments, randomly positioned and oriented within the droplet volume (Fig. 4a and 4e). We note that filament flexibility is fully incorporated into our model, and the filaments are initialized in straight configurations to approximate their equilibrium state in a good solvent. Starting from straight filaments also allows us to isolate and systematically analyze the underlying deposition mechanisms without introducing additional variability arising from arbitrary initial curvature. During evaporation, filaments are first gradually advected toward the contact line (Fig. 4b and 4f). Bending arises primarily from solvation forces at the interface, which cause free chain segments to align with the local curvature of the droplet surface. Additional shape changes result from interparticle interactions and adhesion between deposited segments and the substrate. In this stage, filament segments with sufficient substrate friction overcome solvation forces and remain deposited, while unconstrained segments are carried inward by the receding interface (Fig. 4c and 4g). Filament segments near the contact line get effectively immobilized by friction with the substrate. When one end of a filament is anchored at the contact line, the resulting mechanical constraint is transmitted along the chain via interparticle interactions, influencing segments farther from the contact line and promoting alignment along the inward radial direction towards the droplet center. The filaments finally settle onto the substrate (Fig. 4d and 4h[53], leading to a coffee-ring pattern in the diffusion-limited regime ($K\approx 0$) and a homogeneous deposit in the reaction-limited regime ($K=10$).
@@ III.2.1 Order of Filament Deposits
where $r_{i}$ denotes the radial position of the bin and $N_{i}$ the number of segments within it. Here, a segment is defined as the bond connecting two consecutive monomers within a filament. The vector $\mathbf{u}_{j}$ describes the local orientation of the $j$-th segment, and $\hat{\mathbf{t}}_{j}$ is the local tangent unit vector evaluated at the segment midpoint $\mathbf{r}_{j}=(x_{j},y_{j})$:
@@
Here, $n(r)=N(r)/\Delta A(r)$ is the number density within the $i$-th annular bin $r_{i},r_{i+1}]$, where $N(r)$ is the number of beads located within the bin, and $\Delta A(r)=\pi(r_{i+1}^{2}-r_{i}^{2})$ is the area of the annulus. The average density is ${\color[rgb]{0,0,0}n_{0}}=N_{\text{total}}/(\pi R^{2})$, where $N_{\text{total}}$ is the total number of beads. Values of $g(r)>1$ indicate enrichment, while $g(r)<1$ signifies depletion at the corresponding radial distance. Fig. 5 presents $S_{t}(r)$ and $g(r)$ for four representative cases obtained by varying the filament length $L$ and the filament concentration $c_{v}$. To improve statistical significance, data were first averaged over 16 independent initial configurations for each $(L,c_{v})$ parameter set and subsequently over parameter ranges defined as “small filaments” ($L_{\mathrm{min}}\leq L<L_{\mathrm{small}}$) and “large filaments” ($L_{\mathrm{large}}<L\leq L_{\mathrm{max}}$), as well as “low concentrations” ($c_{\mathrm{min}}\leq c_{v}<c_{\mathrm{low}}$) and “high concentrations” ($c_{\mathrm{high}}<c_{v}\leq c_{\mathrm{max}}$). The corresponding thresholds were $L_{\mathrm{min}}=3$, $L_{\mathrm{small}}=11$, $L_{\mathrm{large}}=23$, $L_{\mathrm{max}}=31$, $c_{\mathrm{min}}=0.0012$, $c_{\mathrm{low}}=0.0035$, $c_{\mathrm{high}}=0.0107$, and $c_{\mathrm{max}}\approx 0.0131$. The averaging was performed with step sizes of $\Delta L=2$ and $\Delta c_{v}=0.0012$. Fig. S1 and S2 in the Supporting Information show results that are averaged only over different initial configurations, but not over different $(L,c_{v})$ parameter values. Within these subsets, no systematic trends are observed. Three characteristic radial regions can be distinguished. As shown in Fig. 5, in the outer region ($\tilde{r}\gtrsim 0.75$), where $\tilde{r}=r/a_{max}$ and $a_{max}=32.79$ denotes the average maximal deposit radius, we find $S_{t}(r)>0$, indicating pronounced tangential alignment (corresponding to Fig. 4d and 4h). This alignment results from outward capillary flows during the CCR phase, with filaments advected to the interface adapting to its curvature. In the intermediate region ($0.25\lesssim\tilde{r}<0.75$), $S_{t}(r)<0$ reflects radial alignment. Here, shear gradients during the CCR phase stretch filaments radially, while in the CCA phase the receding contact line reinforces this orientation as anchored segments resist the inward motion of free segments. In the inner region ($\tilde{r}\lesssim 0.25$), high particle density or low radial flow and excluded-volume effects suppress radial alignment, leading to weak tangential ordering as filaments aggregate into an approximately isotropic central cluster.
@@
The magnitude and spatial extent of these alignment patterns depend on filament length: longer filaments show stronger tangential alignment, particularly outside the inner region, which can be attributed to their lower effective number of inter-particle contacts at the same filament concentration. In contrast, shorter filaments experience more frequent interactions that disrupt orientational order (top row of Fig. 5).
Evaporation conditions further modulate this behavior. In the diffusion-limited regime ($K\approx 0$), strong capillary flows persist into the CCA phase (see Fig. 3), transporting particles toward the contact line and enhancing tangential alignment (see Fig. 5). In contrast, in the reaction-limited regime ($K=10$), uniform evaporation suppresses capillary flows and inward contact line motion dominates, resulting in weaker tangential order. For short filaments, the difference is negligible at low concentrations, where inter-particle interactions are reduced.
The radial density profile $g(r)$ supports these trends. For $K=10$, central enrichment and edge depletion are observed: short filaments exhibit a weak coffee-ring peak near $\tilde{r}\approx 0.85$, whereas long filaments remain concentrated near the center and show no distinct coffee-ring peak. Under $K\approx 0$, central depletion and edge accumulation occur, yet long filaments still exhibit a more homogeneous distribution.